OpenStudy (anonymous):

A garden is rectangular in shape. A row of rose bushes is planted along the diagonal length of 27 feet of the garden. The length of the garden is 18 feet. Which expression can be used to find the width of the garden?

OpenStudy (anonymous):



OpenStudy (anonymous):

then do Pythagorean theorm. 18^2+b^2=27^2.

OpenStudy (anonymous):

rounded: 20

OpenStudy (anonymous):

so would \[\sqrt{27^{2}-18^{2}}\] work?

OpenStudy (anonymous):

no, well, think about it the square root would cancel the squares so you would be left with 27-18. and that gives you 9.

OpenStudy (anonymous):

no you square the numbers then subtract then and then find the square root of that number wich is 20

OpenStudy (anonymous):

oh, ya- sorry. it does work!

OpenStudy (anonymous):

Thanks you for your help though

OpenStudy (anonymous):

np